# Securing the AI Supply Chain: What Can We Learn From Developer-Reported Security Issues and Solutions of AI Projects?

### v1

@@ Abstract.
conference: 48th International Conference on Software Engineering; April 12–18, 2026; Rio De Janeiro, Braziljournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Security and privacy Software and application security
@@ 1. Introduction
With the emergence of Artificial Intelligence (AI)-enabled software, developers of these projects face unique security challenges due to the growing complexity of the AI supply chain. They contend with not only the risks inherited from traditional software, e.g., vulnerable dependencies and insecure coding practices (reichert2024software), but also novel AI-specific threats such as adversarial manipulations (qin2024adversarial). Consequently, multiple high-level risk frameworks, e.g., the OWASP Top 10 for Large Language Models (LLMs) Applications (owasp2025top10), have been developed. These frameworks provide valuable guidance on the theoretical dimensions of AI supply chain risks. However, a significant gap remains between these theoretical frameworks and concrete security problems across different AI supply chain components that developers actually face. Specifically, it is unclear what security issues are commonly encountered, what challenges they pose, and how developers can effectively mitigate or resolve them.
Discussions on open-source platforms such as Hugging Face (HF) and GitHub (GH) can provide insights into real-world security issues in different components of AI-enabled projects. These platforms are central hubs where developers not only share models, code, and datasets but also engage in critical discussions that drive innovation and problem-solving (dabbish2012social; jones2024we). Security-wise, discussions on these platforms include the security of application, source code, and infrastructure on GH (pletea2014security; buhlmann2022developers; horawalavithana2019mentions), as well as the security of AI models, e.g., model serialization vulnerabilities on HF (pepe2024hugging). An example of such security-related discussions is the Ultralytics supply chain attack report (ultralytic_attack), where multiple issues were raised from downstream projects (rten_affect; comfyui_affect; comictranslate_affect) when developers found that the model package was compromised to install crypto miners. This attack demonstrates how a single vulnerability in the AI supply chain can have widespread consequences. The community discussions raised during the attack have proven to be critical data sources in understanding the practical security landscape for developing AI-enabled projects. However, to the best of our knowledge, there is no prior empirical effort to identify and analyze these large-scale discussions to distill real-world security issues that affect the AI supply chain and potential solutions to address these issues.
@@
  • We release our source code, models, datasets as a replication package (replication_package), and our full taxonomy is publicly available at (tax_viz).
@@ 2.1. Security of the AI supply chain
Security challenges of the AI supply chain have become increasingly prominent due to the proliferation of reusable components. While the software supply chain focuses primarily on source code dependencies (reichert2024software), the AI supply chain encompasses not only software components but also unique AI assets such as models, datasets, and specialized development platforms (ding2025rusty). This makes AI systems susceptible not only to traditional vulnerabilities but also to novel attacks, e.g., model poisoning and data manipulations (casey2024large; zhao2024models; jiang2022empirical; tidjon2022threat; qin2024adversarial). All of these problems fall under a broad definition of AI security issues, i.e., any adversarial or accidental conditions that undermine the confidentiality, integrity, availability, and, in turn, the trustworthiness of AI assets and systems (gohil2025managing; pispa2024comprehensive; bieringer2024position). We adopt this definition consistently throughout this study.
Current research on AI supply chain security, while valuable, offers a limited perspective. Studies have used static analysis to identify severe vulnerabilities in models (kathikar2023assessing), highlighted the exploitability of unsafe hosted models (casey2024large), and demonstrated how model reuse can amplify these risks through model artifacts (qin2024adversarial; jiang2023empirical). While essential for finding technical vulnerabilities, these static analysis and artifact-based approaches fail to capture the practical, day-to-day security issues that developers actually encounter, discuss, and resolve. Moreover, analyzing and understanding AI security issues currently depends on high-level security frameworks (weidinger2022taxonomy; cui2024risk; owasp2025top10; mitreatlas), which are derived from theoretical models in the literature. This creates a gap between theoretical risks and real-world developer concerns. This study fills that critical gap by using developer discussions on community platforms as a direct lens into real-world security challenges and solutions of the AI supply chain.
@@ 2.2. Developer discussions on AI security
Open-source platforms, e.g., Hugging Face (HF) and GitHub (GH), are foundational to AI development, serving as registries for pre-trained models, code bases, and problem solving (ding2025rusty; kathikar2023assessing; jiang2023empirical). Issues reported on these platforms offer insight into day-to-day challenges in deploying and maintaining AI systems. For instance, studies of open-source AI repositories show that developers reported runtime errors, dependencies, and implementation bugs (yang2023users; jiang2023empirical; buhlmann2022developers). While many discussions focus on functional faults, they are also the primary forums where critical, real-world security threats first surface. Developers often raise concerns and investigate issues within project repositories (liu2025empirical) before issues are formally submitted to vulnerability databases, e.g., NVD (nvd2025datafeeds). However, disclosure through such databases can take up to 392 days (rodriguez2018analysis). The Ultralytics supply chain attack (hiddenlayer_ultralytics) serves as a compelling example for this scenario. The attack began when a template injection vulnerability was exploited to publish malicious packages that installed hidden crypto miners. The timeline (snyk_ultralytics) shows a rapid, community-driven detection effort on GH, where developers began reporting “PyPI discrepancy” and questioning the integrity of the package (ultralytic_attack). Downstream projects such as ComfyUI and comic-translate also independently raised issues (rten_affect; comfyui_affect; comictranslate_affect) after detecting crypto mining activity in their applications. Developers utilized these threads to diagnose the incident, assess its impact, and share mitigation strategies, with the main issue thread accumulating over 150 interactions (ultralytic_attack). This incident illustrates that developer discussions are more than bug trackers, they function as an essential warning system for security incidents across the AI supply chain, offering timely and community-sourced insights into emerging threats.
@@
Although security-related discussions are critical data sources for security research, their identification is non-trivial. While studies have attempted to detect security reports on GH manually or with keywords (yang2023users; buhlmann2022developers; zahedi2018empirical), recent empirical findings suggest that these approaches significantly underestimate both the volume and diversity of security issues (ghosh2025wasn; zahedi2018empirical). We found that in our dataset, only about 1% of the issues contained explicit security-related keywords. However, when we applied a DL model, the proportion rose to 14% (see Section 4), aligning with findings from other ecosystems (ghosh2025wasn). To overcome these limitations, we leverage DL models and LLMs, which have shown superior performance in software vulnerability prediction (yokoyama2024identifying; ghosh2025wasn; oyetoyan2021improved; waareus2022security), to extract a comprehensive corpus of AI security discussions. This enriched corpus not only captures a broader set of implicitly stated security issues often missed by keyword-based methods, but it also reduces false positives in sampling for subsequent qualitative analysis of security issues and solutions in AI-enabled projects (yu2023security; morrison2018identifying).
@@ 3. Research Questions
RQ1: What is the extent of security reporting in AI-enabled projects? To understand the extent to which security is reported in AI-enabled projects, we first need a reliable method to identify relevant security discussions at scale. Manual review is infeasible, and keyword-based methods produce biased results with limited coverage (zahedi2018empirical). While DL and LLM-based techniques are effective in other ecosystems (e.g., npm (ghosh2025wasn)), they are underexplored in AI contexts. Therefore, RQ1 first investigates how well learning-based techniques identify security-related discussions. Then, we distill the security reporting patterns and the distribution of these reports across different AI applications and model types.
@@ 4.1. Collection of developer discussions
To focus on AI-enabled projects, we began with HF repositories. Its curated ecosystem for AI components (ait2025suitability) provides a more relevant starting point than the broader landscape of GH (kathikar2023assessing). Due to the lack of standardized linking between HF and GH (kathikar2023assessing), we identified their connections as follows: We selected active HF models with at least one “like” and one “download” to ensure community engagement, which yielded 687,653 models as of January 20, 2025. We then parsed the model cards and README files for hyperlinks, resulting in 179,045 models with potential links. Regular expressions were used to extract URLs pointing to GH, yielding 8,972 distinct HF-GH project linkages. The discussions and issues of these linked projects were then collected from both platforms via their APIs. Since not all GH repositories enabled the “discussion” feature, GH discussions and issues were considered as part of the same GH data source. Pull requests and their comments were also included in this source, as the GH API represents both of these as “issues” (github-issues-docs). Conversely, we only collected the discussions from HF, as HF repositories did not have a dedicated “issue” reporting section. We then removed 86,421 discussions and issues raised by 36 bot accounts, identified by filtering for account names with the keyword ‘‘bot’’ and confirming their nature via manual inspection. This resulted in our dataset with 205,667 HF discussions and 2,022,955 GH issues and discussions.
@@ 4.2. Identification of initial security discussions
To select an initial set of AI security discussions, we first curated a list of popular security keywords, including: “vulnerability”, ‘‘vulnerabilities’’, “security”, ‘‘cve’’, ‘‘CVE’’, ‘‘cwe’’ and ‘‘CWE’’. We limited our keywords to only unambiguous security terms and widely used standards (i.e., CVE (cve2024) and CWE (cwe2024)), as we did not want to explicitly target vulnerabilities in a specific area. The keyword search was then applied to all text fields on the collected dataset, e.g., titles and body text, yielding 252 HF Discussions, 207 GH Discussions, and 23,755 GH Issues. However, developers usually reported different artifacts (e.g., built-in logs, chat messages) that contained these keywords, but the discussions themselves were not directly related to security. For example, a developer in openvinotoolkit/openvino-24013 reported a log file that contained the “vulnerabilities” keyword while asking for model serving implementations. Therefore, we manually labeled all the filtered Discussions from HF and GH, but we manually validated only a subset of 264 GH Issues with at least three keywords, as those with fewer keywords were too numerous for exhaustive manual labeling. This threshold allowed us to balance coverage with the available manual labeling effort while still capturing a meaningful set of potentially security-relevant issues. Many of the missing security discussions (e.g., with fewer than three keywords) were expected to be auto-collected using learning-based classifiers in Section 4.3. We used the label “1” (security-related) for problems where the integrity, confidentiality, or availability of AI models, datasets, and associated software components were affected, potentially leading to security vulnerabilities. The label “0” (not related to security) was applied to the remaining records. To minimize subjectivity, the first author independently labeled the whole set and presented the labeling process in frequent meetings with the second and third authors. Disagreements were resolved through discussions and mutual agreements, and the process ended when all labels were assigned.
@@ 4.3. Security discussions classifier development
Training, validation, and testing data preparation. The manually labeled dataset of 252 HF Discussions as well as 471 GH Discussions and Issues in Section 4.2 was shuffled and stratified split by the security label into training and validation subsets with a ratio of 80:20 (nguyen2022vulcurator; yokoyama2024identifying; zhou2021finding; cabrera2021commit2vec; wang2021patchrnn). This step ensured the ratio of security and non-security records remained roughly the same for both sets. To ensure that the testing set was not biased toward the keywords in the training data, we further randomly sampled 100 issues that did not contain keywords from all data sources. After applying the same manual labeling process used while labeling the training/validation sets, we produced a testing set that contained 18 label “1” and 82 label “0” records. We report both datasets and their labels in Table 1. The datasets, hyperparameters, optimal model settings, and prompts are available in our replication package (replication_package).
@@
Deep Learning (DL) models. Fine-tuning DL models have demonstrated significant performance for classifying security-related content (ghosh2025wasn; sharma2024distilbert; liu2024spam). This process adapts a pre-trained model to our security classification task by training it further on our labeled dataset, allowing it to learn the linguistic patterns and security-related contexts. The trained model can then predict whether a discussion is security-related. We utilized three BERT-based models, including the original BERT (devlin2019bert), RoBERTa (liu2019roberta) and distilBERT (sanh2019distilbert), as they have been widely adopted for natural language processing (NLP) and security-related tasks (bharadwaj2022github; minina2023detecting; cao2022sbrpbert; varenov2021security; liu2024spam; sharma2024distilbert; wang2021patchrnn). We also included two fine-tuned variants, secBERT (secBERT) and secureBERT (aghaei2022securebert), as they have achieved better performance in different tasks with cybersecurity corpus (aghaei2022securebert) than BERT and RoBERTa by leveraging domain-specific data and formal security documents. We leveraged the original implementations of these models and only modified the last layer to be used as a binary classification head. The models were then trained with the training/validation set described previously, using the content of the discussions as input text feature, and the security label as the binary classification task label. To ensure reliable training, we performed stratified 10-fold cross-validation on the training/validation set to ensure the security label ratio was roughly the same for each fold, together with early stopping to limit overfitting and randomized grid search to obtain the best hyperparameters for each model. The final performance was measured with the best-performing hyperparameters using the dedicated testing set described earlier.
@@
Large Language Models (LLMs). LLMs can perform classification directly by relying on its pretrained knowledge. By prompting our task in natural language (zero-shot) and adding examples (few-shot)  (brown2020language), LLMs can infer task-specific patterns without additional training, which can be used to predict whether a discussion is security-related. We focused our evaluation on notable open-source LLMs to perform the security classification: llama3 (grattafiori2024llama), phi4 (abdin2024phi), mistral (ollama2023mistral), and deepseek-r1 (guo2025deepseek). llama3 and phi are reasoning-focused models, with phi trained on textbooks and synthetic data. mistral and deepseek aim to balance speed and quality through architectural and training optimizations. We did not include commercial LLMs due to their black-box nature and potential for unannounced updates, which can lead to changed behaviors (chen2024chatgpt). For the selected LLMs, we employed zero-shot and few-shot prompting (brown2020language) with a temperature of 0 to obtain deterministic output, and evaluated these models using the testing set described earlier. In the few-shot setting, we included six examples from our training data, three of which were security-centric to provide sufficient context without overwhelming the prompts and mitigate risks of irrelevant output (sheng2025llms). The prompts were then customized according to the reference from (yokoyama2024identifying).
@@
Evaluation metrics. Matthews Correlation Coefficient (MCC) and Macro F1-Score were used to quantify the performance of the classifiers. F1-Score, the harmonic mean of precision and recall, balances false positives and false negatives with values ranging from 0 to 1 (best). MCC, calculated from all four values of the confusion matrix, ensures a more balance metric and ranges from -1 to 1 (best). Since MCC can better reflect performance on imbalanced datasets (chicco2020advantages), we reported both but used MCC as the primary indicator.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3787796isbn: 979-8-4007-2025-3/2026/04ccs: Security and privacy Software and application security
@@ 1. Introduction
With the emergence of Artificial Intelligence (AI)-enabled software, developers of these projects face unique security challenges due to the growing complexity of the AI supply chain. They contend with not only the risks inherited from traditional software, e.g., vulnerable dependencies and insecure coding practices (Reichert and Obelheiro, 2024), but also novel AI-specific threats such as adversarial manipulations (Qin et al., 2024). Consequently, multiple high-level risk frameworks, e.g., the OWASP Top 10 for Large Language Models (LLMs) Applications (Foundation, 2025), have been developed. These frameworks provide valuable guidance on the theoretical dimensions of AI supply chain risks. However, a significant gap remains between these theoretical frameworks and concrete security problems across different AI supply chain components that developers actually face. Specifically, it is unclear what security issues are commonly encountered, what challenges they pose, and how developers can effectively mitigate or resolve them.
Discussions on open-source platforms such as Hugging Face (HF) and GitHub (GH) can provide insights into real-world security issues in different components of AI-enabled projects. These platforms are central hubs where developers not only share models, code, and datasets but also engage in critical discussions that drive innovation and problem-solving (Dabbish et al., 2012; Jones et al., 2024). Security-wise, discussions on these platforms include the security of application, source code, and infrastructure on GH (Pletea et al., 2014; Bühlmann and Ghafari, 2022; Horawalavithana et al., 2019), as well as the security of AI models, e.g., model serialization vulnerabilities on HF (Pepe et al., 2024). An example of such security-related discussions is the Ultralytics supply chain attack report (Ultralytics, ), where multiple issues were raised from downstream projects (robertknight, ; ltdrdata, ; ogkalu2, ) when developers found that the model package was compromised to install crypto miners. This attack demonstrates how a single vulnerability in the AI supply chain can have widespread consequences. The community discussions raised during the attack have proven to be critical data sources in understanding the practical security landscape for developing AI-enabled projects. However, to the best of our knowledge, there is no prior empirical effort to identify and analyze these large-scale discussions to distill real-world security issues that affect the AI supply chain and potential solutions to address these issues.
@@
  • We release our source code, models, datasets as a replication package (Nguyen, 2025b), and our full taxonomy is publicly available at (Nguyen, 2025a).
@@ 2.1. Security of the AI supply chain
Security challenges of the AI supply chain have become increasingly prominent due to the proliferation of reusable components. While the software supply chain focuses primarily on source code dependencies (Reichert and Obelheiro, 2024), the AI supply chain encompasses not only software components but also unique AI assets such as models, datasets, and specialized development platforms (Ding et al., 2025). This makes AI systems susceptible not only to traditional vulnerabilities but also to novel attacks, e.g., model poisoning and data manipulations (Casey et al., 2024; Zhao et al., 2024a; Jiang et al., 2022; Tidjon and Khomh, 2022; Qin et al., 2024). All of these problems fall under a broad definition of AI security issues, i.e., any adversarial or accidental conditions that undermine the confidentiality, integrity, availability, and, in turn, the trustworthiness of AI assets and systems (Gohil, 2025; Pispa and Halunen, 2024; Bieringer et al., 2024). We adopt this definition consistently throughout this study.
Current research on AI supply chain security, while valuable, offers a limited perspective. Studies have used static analysis to identify severe vulnerabilities in models (Kathikar et al., 2023), highlighted the exploitability of unsafe hosted models (Casey et al., 2024), and demonstrated how model reuse can amplify these risks through model artifacts (Qin et al., 2024; Jiang et al., 2023). While essential for finding technical vulnerabilities, these static analysis and artifact-based approaches fail to capture the practical, day-to-day security issues that developers actually encounter, discuss, and resolve. Moreover, analyzing and understanding AI security issues currently depends on high-level security frameworks (Weidinger et al., 2022; Cui et al., 2024; Foundation, 2025; Corporation, 2025), which are derived from theoretical models in the literature. This creates a gap between theoretical risks and real-world developer concerns. This study fills that critical gap by using developer discussions on community platforms as a direct lens into real-world security challenges and solutions of the AI supply chain.
@@ 2.2. Developer discussions on AI security
Open-source platforms, e.g., Hugging Face (HF) and GitHub (GH), are foundational to AI development, serving as registries for pre-trained models, code bases, and problem solving (Ding et al., 2025; Kathikar et al., 2023; Jiang et al., 2023). Issues reported on these platforms offer insight into day-to-day challenges in deploying and maintaining AI systems. For instance, studies of open-source AI repositories show that developers reported runtime errors, dependencies, and implementation bugs (Yang et al., 2023; Jiang et al., 2023; Bühlmann and Ghafari, 2022). While many discussions focus on functional faults, they are also the primary forums where critical, real-world security threats first surface. Developers often raise concerns and investigate issues within project repositories (Liu et al., 2025) before issues are formally submitted to vulnerability databases, e.g., NVD (of Standards and (NIST), 2025). However, disclosure through such databases can take up to 392 days (Rodriguez et al., 2018). The Ultralytics supply chain attack (Layer, 2023) serves as a compelling example for this scenario. The attack began when a template injection vulnerability was exploited to publish malicious packages that installed hidden crypto miners. The timeline (Team, 2023) shows a rapid, community-driven detection effort on GH, where developers began reporting “PyPI discrepancy” and questioning the integrity of the package (Ultralytics, ). Downstream projects such as ComfyUI and comic-translate also independently raised issues (robertknight, ; ltdrdata, ; ogkalu2, ) after detecting crypto mining activity in their applications. Developers utilized these threads to diagnose the incident, assess its impact, and share mitigation strategies, with the main issue thread accumulating over 150 interactions (Ultralytics, ). This incident illustrates that developer discussions are more than bug trackers, they function as an essential warning system for security incidents across the AI supply chain, offering timely and community-sourced insights into emerging threats.
@@
Although security-related discussions are critical data sources for security research, their identification is non-trivial. While studies have attempted to detect security reports on GH manually or with keywords (Yang et al., 2023; Bühlmann and Ghafari, 2022; Zahedi et al., 2018; Le et al., 2020), recent empirical findings suggest that these approaches significantly underestimate both the volume and diversity of security issues (Ghosh et al., 2025; Zahedi et al., 2018). We found that in our dataset, only about 1% of the issues contained explicit security-related keywords. However, when we applied a DL model, the proportion rose to 14% (see Section 4), aligning with findings from other ecosystems (Ghosh et al., 2025). To overcome these limitations, we leverage DL models and LLMs, which have shown superior performance in software vulnerability prediction (Yokoyama et al., 2024; Le et al., 2024; Ghosh et al., 2025), to extract a comprehensive corpus of AI security discussions. This enriched corpus not only captures a broader set of implicitly stated security issues often missed by keyword-based methods, but it also reduces false positives in sampling for subsequent qualitative analysis of security issues and solutions in AI-enabled projects (Yu et al., 2023; Morrison et al., 2018).
@@ 3. Research Questions
RQ1: What is the extent of security reporting in AI-enabled projects? To understand the extent to which security is reported in AI-enabled projects, we first need a reliable method to identify relevant security discussions at scale. Manual review is infeasible, and keyword-based methods produce biased results with limited coverage (Zahedi et al., 2018). While DL and LLM-based techniques are effective in other ecosystems (e.g., npm (Ghosh et al., 2025)), they are underexplored in AI contexts. Therefore, RQ1 first investigates how well learning-based techniques identify security-related discussions. Then, we distill the security reporting patterns and the distribution of these reports across different AI applications and model types.
@@ 4.1. Collection of developer discussions
To focus on AI-enabled projects, we began with HF repositories. Its curated ecosystem for AI components (Ait et al., 2025) provides a more relevant starting point than the broader landscape of GH (Kathikar et al., 2023). Due to the lack of standardized linking between HF and GH (Kathikar et al., 2023), we identified their connections as follows: We selected active HF models with at least one “like” and one “download” to ensure community engagement, which yielded 687,653 models as of January 20, 2025. We then parsed the model cards and README files for hyperlinks, resulting in 179,045 models with potential links. Regular expressions were used to extract URLs pointing to GH, yielding 8,972 distinct HF-GH project linkages. The discussions and issues of these linked projects were then collected from both platforms via their APIs. Since not all GH repositories enabled the “discussion” feature, GH discussions and issues were considered as part of the same GH data source. Pull requests and their comments were also included in this source, as the GH API represents both of these as “issues” (Docs, 2022). Conversely, we only collected the discussions from HF, as HF repositories did not have a dedicated “issue” reporting section. We then removed 86,421 discussions and issues raised by 36 bot accounts, identified by filtering for account names with the keyword ‘‘bot’’ and confirming their nature via manual inspection. This resulted in our dataset with 205,667 HF discussions and 2,022,955 GH issues and discussions.
@@ 4.2. Identification of initial security discussions
To select an initial set of AI security discussions, we first curated a list of popular security keywords, including: “vulnerability”, ‘‘vulnerabilities’’, “security”, ‘‘cve’’, ‘‘CVE’’, ‘‘cwe’’ and ‘‘CWE’’. We limited our keywords to only unambiguous security terms and widely used standards (i.e., CVE (of Standards and (NIST), 2024) and CWE (Corporation, 2024)), as we did not want to explicitly target vulnerabilities in a specific area. The keyword search was then applied to all text fields on the collected dataset, e.g., titles and body text, yielding 252 HF Discussions, 207 GH Discussions, and 23,755 GH Issues. However, developers usually reported different artifacts (e.g., built-in logs, chat messages) that contained these keywords, but the discussions themselves were not directly related to security. For example, a developer in openvinotoolkit/openvino-24013 reported a log file that contained the “vulnerabilities” keyword while asking for model serving implementations. Therefore, we manually labeled all the filtered Discussions from HF and GH, but we manually validated only a subset of 264 GH Issues with at least three keywords, as those with fewer keywords were too numerous for exhaustive manual labeling. This threshold allowed us to balance coverage with the available manual labeling effort while still capturing a meaningful set of potentially security-relevant issues. Many of the missing security discussions (e.g., with fewer than three keywords) were expected to be auto-collected using learning-based classifiers in Section 4.3. We used the label “1” (security-related) for problems where the integrity, confidentiality, or availability of AI models, datasets, and associated software components were affected, potentially leading to security vulnerabilities. The label “0” (not related to security) was applied to the remaining records. To minimize subjectivity, the first author independently labeled the whole set and presented the labeling process in frequent meetings with the second and third authors. Disagreements were resolved through discussions and mutual agreements, and the process ended when all labels were assigned.
@@ 4.3. Security discussions classifier development
Training, validation, and testing data preparation. The manually labeled dataset of 252 HF Discussions as well as 471 GH Discussions and Issues in Section 4.2 was shuffled and stratified split by the security label into training and validation subsets with a ratio of 80:20 (Nguyen et al., 2022; Yokoyama et al., 2024; Zhou et al., 2021; Cabrera Lozoya et al., 2021; Wang et al., 2021). This step ensured the ratio of security and non-security records remained roughly the same for both sets. To ensure that the testing set was not biased toward the keywords in the training data, we further randomly sampled 100 issues that did not contain keywords from all data sources. After applying the same manual labeling process used while labeling the training/validation sets, we produced a testing set that contained 18 label “1” and 82 label “0” records. We report both datasets and their labels in Table 1. The datasets, hyperparameters, optimal model settings, and prompts are available in our replication package (Nguyen, 2025b).
@@
Deep Learning (DL) models. Fine-tuning DL models have demonstrated significant performance for classifying security-related content (Ghosh et al., 2025; Sharma, 2024; Liu et al., 2024). This process adapts a pre-trained model to our security classification task by training it further on our labeled dataset, allowing it to learn the linguistic patterns and security-related contexts. The trained model can then predict whether a discussion is security-related. We utilized three BERT-based models, including the original BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019) and distilBERT (Sanh et al., 2019), as they have been widely adopted for natural language processing (NLP) and security-related tasks (Bharadwaj and Kadam, 2022; Minina et al., 2023; Cao et al., 2022; Varenov and Gabdrahmanov, 2021; Sharma, 2024; Wang et al., 2021). We also included two fine-tuned variants, secBERT (jackaduma, 2024) and secureBERT (Aghaei et al., 2022), as they have achieved better performance in different tasks with cybersecurity corpus (Aghaei et al., 2022) than BERT and RoBERTa by leveraging domain-specific data and formal security documents. We leveraged the original implementations of these models and only modified the last layer to be used as a binary classification head. The models were then trained with the training/validation set described previously, using the content of the discussions as input text feature, and the security label as the binary classification task label. To ensure reliable training, we performed stratified 10-fold cross-validation on the training/validation set to ensure the security label ratio was roughly the same for each fold, together with early stopping to limit overfitting and randomized grid search to obtain the best hyperparameters for each model. The final performance was measured with the best-performing hyperparameters using the dedicated testing set described earlier.
@@
Large Language Models (LLMs). LLMs can perform classification directly by relying on its pretrained knowledge. By prompting our task in natural language (zero-shot) and adding examples (few-shot)  (Brown et al., 2020), LLMs can infer task-specific patterns without additional training, which can be used to predict whether a discussion is security-related. We focused our evaluation on notable open-source LLMs to perform the security classification: llama3 (Grattafiori et al., 2024), phi4 (Abdin et al., 2024), mistral (Ollama, 2023), and deepseek-r1 (Guo et al., 2025). llama3 and phi are reasoning-focused models, with phi trained on textbooks and synthetic data. mistral and deepseek aim to balance speed and quality through architectural and training optimizations. We did not include commercial LLMs due to their black-box nature and potential for unannounced updates, which can lead to changed behaviors (Chen et al., 2024). For the selected LLMs, we employed zero-shot and few-shot prompting (Brown et al., 2020) with a temperature of 0 to obtain deterministic output, and evaluated these models using the testing set described earlier. In the few-shot setting, we included six examples from our training data, three of which were security-centric to provide sufficient context without overwhelming the prompts and mitigate risks of irrelevant output (Sheng et al., 2025). The prompts were then customized according to the reference from (Yokoyama et al., 2024).
@@
Evaluation metrics. Matthews Correlation Coefficient (MCC) and Macro F1-Score were used to quantify the performance of the classifiers. F1-Score, the harmonic mean of precision and recall, balances false positives and false negatives with values ranging from 0 to 1 (best). MCC, calculated from all four values of the confusion matrix, ensures a more balanced metric and ranges from -1 to 1 (best). MCC can better reflect performance on imbalanced datasets (Nguyen et al., 2024; Le and Ali Babar, 2024; Le and Babar, 2024), so we reported both but used MCC as the primary indicator.